# Federated Granger Causality Learning for Interdependent Clients with State Space Representation

### v1

@@ 3 Problem Setting
A discussion on the possible choices of $f_{c}$, $f_{a}$, $f_{ML}$ along with the rationale behind our models, is provided in the Appendix A.2. A simplified pictorial description of the aforementioned problem setting is shown in figure 1. A pseudocode for our proposed framework is given in Appendix A.3.
@@ 4 Federated Granger Causality Framework
The client model $f_{c}(.)$ is a Kalman filter with access to client-specific measurements $y_{m}\forall m\in\{1,...,M\}$. It uses only the diagonal blocks of the state matrix $A$ and output matrix $C$ (given by $A_{mm}$, and $C_{mm}$ respectively) . Eqs in the first column of Table 1 define the client model, using only $A_{mm}$, and $C_{mm}$ which are known apriori. If unknown, they can be estimated locally using $y_{m}$. The estimated and predicted states are ${(\hat{h}^{t-1}_{m})}_{c}$, and ${(h^{t}_{m})}_{c}$, with residual ${(r^{t}_{m})}_{c}$ and Kalman gain ${(K_{m})}_{c}$.
@@ 4.1 Augmented Client Model
$=\theta_{m}^{k}-\eta_{1}\cdot\nabla_{\theta_{m}^{k}}{(L_{m})}_{a}-\eta_{2}\cdot\big{[}\nabla_{{(\hat{h}_{m}^{t})}_{a}}{L}_{s}\cdot\nabla_{\theta_{m}^{k}}{{(\hat{h}_{m}^{t})}_{a}}\big{]}$ (2)
@@ 5 Understanding Decentralization Through a Centralized Lens
Next, we state proposition 5.3 that gives the values of the optimal augmented client and server model parameters given by $\theta_{m}^{*}$ and $\hat{A}_{mn}^{*}$ respectively. While the first condition gives the closed-form for $\theta_{m}^{*}$ as a function of the knowns; the second condition gives $\hat{A}_{mn}^{*}$ as a function of $\theta_{m}^{*}$.
@@
If $A_{mm}\neq 0$, when augmented client and server model parameter converges to $\theta_{m}^{*}$ and $\hat{A}_{mn}^{*},n\neq m$, respectively then the following holds:
@@
    $A_{mm}\theta_{m}^{*}y^{t}_{m}-\sum_{n\neq m}\hat{A}^{*}_{mn}{(\hat{h}_{n}^{t})}_{c}=0$
@@
The federated framework effectively solves the following recurrent equation:
@@
$J$ $:=\begin{bmatrix}0&0&\ldots&\mathrm{vec}(D)\end{bmatrix}^{T}\hskip 2.84544pt\text{with}\hskip 2.84544ptD:=A_{mm}^{T}C_{mm}^{T}{(r^{t-1}_{m})}_{c}{y^{t-1}_{m}}^{T}\hskip 2.84544pt\text{and}\hskip 2.84544pt\otimes\hskip 2.84544pt\text{is the Kronecker prod.}$.
@@ 6 Asymptotic Convergence to the Centralized Oracle
Let ${(\hat{h}_{m}^{t,k})}_{a}:={(\hat{h}_{m}^{t})}_{c}+\theta_{m}^{k}y^{t}_{m}$, and ${({h}_{m}^{t,k})}_{a}:=A_{mm}\cdot{(\hat{h}_{m}^{t,k})}_{a}$ (see table 1). Then, the following convergence result holds: $\lim_{k\to\infty}\mathbb{E}[\lVert{(h_{m}^{t,k})}_{a}-{(h_{m}^{t})}_{o}\rVert=0\hskip 5.69046pt\forall m\in\{1,...,M\}$
@@ A.4 Proofs
$H_{s}^{t}=\begin{pmatrix}{(h_{1}^{t})}_{s}\\ .\\ .\\ .\\ {(h_{m}^{t})}_{s}\end{pmatrix}=\begin{pmatrix}A_{11}{(\hat{h}^{t-1}_{1})}_{c}+\sum_{n\neq 1}\hat{A}_{1n}^{k}{(\hat{h}^{t-1}_{n})}_{c}\\ .\\ .\\ .\\ A_{mm}{(\hat{h}^{t-1}_{m})}_{c}+\sum_{n\neq m}\hat{A}_{mn}^{k}{(\hat{h}^{t-1}_{n})}_{c}\end{pmatrix}$ (14)
@@
$\begin{split}\hat{A}_{mn}^{k+1}=\hat{A}_{mn}^{k}+2\gamma\bigg{(}A_{mm}(\hat{h}^{t-1}_{m})_{a}-\big{[}A_{mm}(\hat{h}^{t-1}_{m})_{c}+\sum_{n\neq m}\hat{A}_{mn}^{k}(\hat{h}^{t-1}_{n})_{c}\big{]}\bigg{)}{(\hat{h}^{t-1}_{n})}_{c}^{T}\end{split}$ (17)
@@
A.4.3 Proof of Proposition 5.3:
Condition (1): Convergence of $\theta^{k}_{m}$ implies $\nabla_{\theta^{k}_{m}}{(L_{m})}_{a}=0$. Therefore using equation 21 we get,
$y_{m}^{t}-C_{mm}\cdot A_{mm}\cdot\big{[}(\hat{h}^{t-1}_{m})_{c}+\theta_{m}^{*}y_{m}^{t-1}\big{]}=0$ (25)
Condition (2): Convergence of $\hat{A}^{k}_{mn}$ implies $\nabla_{\hat{A}^{k}_{mn}}L_{s}=0$. Using equation 16 we obtain,
$A_{mm}(\hat{h}^{t-1}_{m})_{a}-\big{[}A_{mm}(\hat{h}^{t-1}_{m})_{c}+\sum_{n\neq m}\hat{A}_{mn}^{*}(\hat{h}^{t-1}_{n})_{c}\big{]}=0$ (26)
We also know the following from the definition of augmented client model:
${(\hat{h}^{t-1}_{m})}_{a}={(\hat{h}^{t-1}_{m})}_{c}+\theta_{m}^{*}y_{m}^{t-1}$ (27)
Substituting equation 27 in equation 26 we have,
$A_{mm}\theta_{m}^{*}y_{m}^{t-1}-\sum_{n\neq m}\hat{A}_{mn}^{*}(\hat{h}^{t-1}_{n})_{c}=0$ (28)
@@
To handle the matrix equations 29 and 30, we first linearize them using the vectorization technique described in definitions A.1.
@@
We observe many terms in equations 31 and 32 contain multiplication of two or three matrices. To vectorize such terms we utilize definition A.2
@@
$J$ $:=\begin{bmatrix}0&0&\ldots&\mathrm{vec}(D)\end{bmatrix}^{T}\hskip 2.84544pt\text{with}\hskip 2.84544ptD:=A_{mm}^{T}C_{mm}^{T}{(r^{t-1}_{m})}_{c}{y^{t-1}_{m}}^{T}\hskip 2.84544pt\text{and}\hskip 2.84544pt\otimes\hskip 2.84544pt\text{is the Kronecker prod.}$.
@@
$\Delta^{*}=H\cdot\Delta*+J$
@@
Substituting $\eta\leq\frac{2}{\mu+\mathcal{L}}$ in inequality 37 we obtain:
@@
If $\rho(A-AK_{o}C)<1$ then the oracle’s expected steady state error is zero and we term such an oracle as convergent. For a convergent oracle the expected residuals $\mathbb{E}[\|r_{o}\rVert]=0$.
@@
Therefore, subtracting both the equations and taking expectation of the $l_{2}$ norm of their difference we obtain:
$\mathbb{E}\big{[}\lVert{(r_{m}^{t})}_{a}-{(r_{m}^{t})}_{o}\rVert\big{]}=\mathbb{E}\big{[}\lVert C_{mm}\cdot({(h_{m}^{t})}_{a}-{(h_{m}^{t})}_{o})\rVert\big{]}$ (40)
We know from definition A.3 that $\mathbb{E}[\lVert{(r_{m}^{t})}_{o}\rVert]=0$.
We also know that $\mathbb{E}[\lVert{(r_{m}^{t})}_{a}\rVert]=0$.
Therefore, $\mathbb{E}\big{[}\lVert({(h_{m}^{t})}_{a}-{(h_{m}^{t})}_{o})\rVert\big{]}=0$
@@
We know that $\mathbb{E}[\lVert{(r_{m}^{t})}_{o}\rVert]=0$. Given $\rho(A_{mm}-A_{mm}{(K_{m})}_{c}C_{mm}<1$, the quantity $\mathbb{E}[\lVert{(r_{m}^{t})}_{c}\rVert]$ is finite.
Therefore from inequality 45, we can say that $\mathbb{E}[\lVert{(\hat{h}_{m}^{t})}_{o}-{(\hat{h}_{m}^{t})}_{c}\rVert]\leq\delta_{max}^{m}$ where, $\delta_{max}^{m}$ is a function of $\mathbb{E}[\lVert{(r_{m}^{t})}_{c}\rVert],C_{mm},A_{mm},\&\sum_{n\neq m}A_{mn}{(\hat{h}_{n}^{t})}_{o}$.
@@
${(r_{m}^{t})^{*}}_{a}={y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{c}+A_{mm}\theta_{m}^{*}y_{m}^{t}\bigg{)}$ (46)
We also know from $\nabla_{\hat{A}_{mn}}L_{s}=0$ the following:
$A_{mm}\theta_{m}^{*}y_{m}^{t}-\sum_{n\neq m}\hat{A}^{*}_{mn}{(\hat{h}_{n}^{t})}_{c}=0$ (47)
@@
${{(r_{m}^{t})}^{*}}_{a}={y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{c}+\sum_{n\neq m}\hat{A}^{*}_{mn}{(\hat{h}_{n}^{t})}_{c}\bigg{)}$ (48)
@@
${(r_{m}^{t})}_{o}={y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{o}+\sum_{n\neq m}{A}_{mn}{(\hat{h}_{n}^{t})}_{o}\bigg{)}$ (49)
Subtracting equation 49 from 48, and using results from proposition 6.2 i.e., $\mathbb{E}[\lVert{(\hat{h}_{m}^{t})}_{o}-{(\hat{h}_{m}^{t})}_{c}\rVert]\leq\delta_{max}^{m}$ we obtain:
@@
Let $\sigma_{min}^{n}=\min_{n\neq m}\mathbb{E}[\lVert{(\hat{h}_{n}^{t-1})}_{o}\rVert]$, then using results of theorem 6.3 and inequality 51 we can infer that, $\big{\lVert}\sum_{n\neq m}[\hat{A}_{mn}^{*}-A_{mn}]\big{\rVert}\leq\frac{1}{\sigma_{min}^{n}}\cdot\bigg{(}\lVert A_{mm}\delta^{m}_{max}\rVert+\lVert\sum_{n,n\neq m}\hat{A}^{*}_{mn}\delta^{n}_{max}\rVert\bigg{)}$

### v4

@@ 3 Problem Setting
A discussion on the possible choices of $f_{c}$, $f_{a}$, $f_{ML}$ along with the rationale behind our models, is provided in the Appendix A.2. A simplified pictorial description of the aforementioned problem setting is shown in figure 1. A pseudocode for our proposed framework is given in Appendix A.3. Readers can find the code of this paper and associated experiments in https://github.com/federated-interdependency-learning/fed_granger_causality.git
@@ 4 Federated Granger Causality Framework
The client model $f_{c}(.)$ is a Kalman filter with access to client-specific measurements $y_{m}\forall m\in\{1,...,M\}$. It uses only the diagonal blocks of the state matrix $A$ and output matrix $C$ (given by $A_{mm}$, and $C_{mm}$ respectively). Equations in the first column of Table 1 define the client model, using only $A_{mm}$, and $C_{mm}$ which are known apriori. If unknown, they can be estimated locally using $y_{m}$. The estimated and predicted states are ${(\hat{h}^{t-1}_{m})}_{c}$, and ${(h^{t}_{m})}_{c}$, with residual ${(r^{t}_{m})}_{c}$ and Kalman gain ${(K_{m})}_{c}$.
@@ 4.1 Augmented Client Model
$=\theta_{m}^{k}-\eta_{1}\cdot\nabla_{\theta_{m}^{k}}{(L_{m})}_{a}-\eta_{2}\cdot\big{[}\nabla_{{(\hat{h}_{m}^{t})}_{a}}{L}_{s}\cdot{y^{t}_{m}}^{\top}\big{]}$ (2)
@@ 5 Understanding Decentralization Through a Centralized Lens
Next, we state proposition 5.3 that gives the values of the optimal augmented client and server model parameters given by $\theta_{m}^{*}$ and $\hat{A}_{mn}^{*}$, respectively. While the first condition gives the closed-form for $\theta_{m}^{*}$ as a function of the knowns; the second condition gives $\hat{A}_{mn}^{*}$ as a function of $\theta_{m}^{*}$.
@@
If $C_{mm}A_{mm}$ is of full rank and $y_{m}^{t}\neq 0\forall t$, when augmented client and server model parameter converges to $\theta_{m}^{*}$ and $\hat{A}_{mn}^{*},n\neq m$, respectively then:
@@
    $\mathbb{E}\big{[}\left(A_{mm}\theta_{m}^{*}y_{m}^{t-1}\right)^{\top}(\hat{h}_{n}^{t-1})_{c}-\sum_{p\neq m}\hat{A}_{mp}^{*}(\hat{h}_{p}^{t-1})_{c}^{\top}(\hat{h}_{n}^{t-1})_{c}\big{]}=0\quad\forall n\neq m$
@@
For any client $m$, the federated framework effectively solves the following recurrent equation:
@@
$P_{mm}$ $:=(I-2\gamma{(\hat{h}^{t-1}_{m})}_{c}{(\hat{h}^{t-1}_{m})}_{c}^{T})\otimes I\hskip 5.69046pt;\hskip 5.69046ptQ_{mn}:=y^{t-1}_{m}{(\hat{h}^{t-1}_{n})}_{c}^{T}\hskip 5.69046pt;\hskip 5.69046ptR:=2A_{mm}\hskip 5.69046pt;\hskip 5.69046ptG:=y^{t-1}_{m}{y^{t-1}_{m}}^{T}$
@@ 6 Asymptotic Convergence to the Centralized Oracle
[…]
@@ 10 Conclusion and Limitations
Acknowledgments
This effort is supported by NASA under grant number 80NSSC19K1052 as part of the NASA Space Technology Research Institute (STRI) Habitats Optimized for Missions of Exploration (HOME) ‘SmartHab’ Project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Aeronautics and Space Administration.
@@ A.4 Proofs
$H_{s}^{t}=\begin{pmatrix}{(h_{1}^{t})}_{s}\\ .\\ .\\ .\\ {(h_{M}^{t})}_{s}\end{pmatrix}=\begin{pmatrix}A_{11}{(\hat{h}^{t-1}_{1})}_{c}+\sum_{p\neq 1}\hat{A}_{1p}^{k}{(\hat{h}^{t-1}_{p})}_{c}\\ .\\ .\\ .\\ A_{MM}{(\hat{h}^{t-1}_{M})}_{c}+\sum_{p\neq M}\hat{A}_{Mp}^{k}{(\hat{h}^{t-1}_{p})}_{c}\end{pmatrix}$ (14)
@@
$\begin{split}\hat{A}_{mn}^{k+1}=\hat{A}_{mn}^{k}+2\gamma\bigg{(}A_{mm}(\hat{h}^{t-1}_{m})_{a}-\big{[}A_{mm}(\hat{h}^{t-1}_{m})_{c}+\sum_{p\neq m}\hat{A}_{mp}^{k}(\hat{h}^{t-1}_{p})_{c}\big{]}\bigg{)}{(\hat{h}^{t-1}_{n})}_{c}^{T}\end{split}$ (17)
@@
A.4.3 Proof of Proposition 5.3
Condition (1):
At convergence of $\theta_{m}^{k}\to\theta_{m}^{*}$, the gradient of the client loss $\nabla_{\theta_{m}}(L_{m})_{a}$ vanishes in expectation. From Equation 21:
$\mathbb{E}\left[\nabla_{\theta_{m}}(L_{m})_{a}\right]=-2\mathbb{E}\left[\big{(}C_{mm}A_{mm}\big{)}^{\top}\left(y_{m}^{t}-C_{mm}A_{mm}\big{[}(\hat{h}^{t-1}_{m})_{c}+\theta_{m}^{*}y_{m}^{t-1}\big{]}\right){y_{m}^{t-1}}^{\top}\right]=0.$ (25)
Assuming $C_{mm}A_{mm}$ is full rank and $y_{m}^{t-1}\neq 0$, this simplifies to:
$\mathbb{E}\left[y_{m}^{t}\right]=C_{mm}A_{mm}\left(\mathbb{E}\left[(\hat{h}^{t-1}_{m})_{c}\right]+\theta_{m}^{*}\mathbb{E}\left[y_{m}^{t-1}\right]\right).$ (26)
This ensures the client’s augmented model reconstructs $\mathbb{E}[y_{m}^{t}]$ unbiasedly (i.e., $\mathbb{E}[{(r_{m}^{t})}_{a}]=0$).
Condition (2): At convergence of $\hat{A}_{mn}^{k}\to\hat{A}_{mn}^{*}$, the gradient of the server loss $\nabla_{\hat{A}_{mn}}L_{s}$ vanishes in expectation. From Equation 16:
$\mathbb{E}\left[\nabla_{\hat{A}_{mn}}L_{s}\right]=-2\mathbb{E}\left[\left(A_{mm}(\hat{h}_{m}^{t-1})_{a}-\big{[}A_{mm}(\hat{h}_{m}^{t-1})_{c}+\sum_{p\neq m}\hat{A}_{mp}^{*}(\hat{h}_{p}^{t-1})_{c}\big{]}\right)(\hat{h}_{n}^{t-1})_{c}^{\top}\right]=0.$ (27)
Substitute $(\hat{h}_{m}^{t-1})_{a}=(\hat{h}_{m}^{t-1})_{c}+\theta_{m}^{*}y_{m}^{t-1}$ (from Equation 18):
$-2\mathbb{E}\left[\left(A_{mm}\theta_{m}^{*}y_{m}^{t-1}-\sum_{p\neq m}\hat{A}_{mp}^{*}(\hat{h}_{p}^{t-1})_{c}\right)(\hat{h}_{n}^{t-1})_{c}^{\top}\right]=0$ (28)
$\mathbb{E}\left[\left(A_{mm}\theta_{m}^{*}y_{m}^{t-1}\right)^{\top}(\hat{h}_{n}^{t-1})_{c}\right]=\sum_{p\neq m}\hat{A}_{mp}^{*}\mathbb{E}\left[(\hat{h}_{p}^{t-1})_{c}^{\top}(\hat{h}_{n}^{t-1})_{c}\right]\quad\forall n\neq m.$ (29)
@@
To handle the matrix equations 30 and 31, we first linearize them using the vectorization technique described in definitions A.1.
@@
We observe many terms in equations 32 and 33 contain multiplication of two or three matrices. To vectorize such terms we utilize definition A.2
@@
$P_{mm}$ $:=(I-2\gamma{(\hat{h}^{t-1}_{m})}_{c}{(\hat{h}^{t-1}_{m})}_{c}^{T})\otimes I\hskip 5.69046pt;\hskip 5.69046ptQ_{mn}:=y^{t-1}_{m}{(\hat{h}^{t-1}_{n})}_{c}^{T}\hskip 5.69046pt;\hskip 5.69046ptR:=2A_{mm}\hskip 5.69046pt;\hskip 5.69046ptG:=y^{t-1}_{m}{y^{t-1}_{m}}^{T}$
@@
$\Delta^{*}=H\cdot\Delta^{*}+J$
@@
Substituting $\eta\leq\frac{2}{\mu+\mathcal{L}}$ in inequality 38 we obtain:
@@
If $\rho(A-AK_{o}C)<1$ then the oracle’s expected steady state error is zero and we term such an oracle as convergent. For a convergent oracle the expected residuals $\mathbb{E}[r_{o}]=0$.
@@
Therefore, subtracting both the equations and taking expectation of their difference we obtain:
$\mathbb{E}\big{[}{(r_{m}^{t})}_{a}-{(r_{m}^{t})}_{o}\big{]}=\mathbb{E}\big{[}C_{mm}\cdot({(h_{m}^{t})}_{a}-{(h_{m}^{t})}_{o})\big{]}$ (41)
We know from definition A.3 that $\mathbb{E}[{(r_{m}^{t})}_{o}]=0$.
We also know that ${(r_{m}^{t})}_{a}=y_{m}^{t}-C_{mm}A_{mm}\left[(\hat{h}^{t-1}_{m})_{c}+\theta_{m}^{*}y_{m}^{t-1}\right]$. Thus, from condition (1) of Proposition 5.3, we know that $\mathbb{E}[{(r_{m}^{t})}_{a}]=0$.
Therefore, using equation 41 we know that, $\mathbb{E}\big{[}C_{mm}({(h_{m}^{t})}_{a}-{(h_{m}^{t})}_{o})\big{]}=0$
If $C_{mm}$ is of full-rank (given) then, $\mathbb{E}\big{[}({(h_{m}^{t})}_{a}-{(h_{m}^{t})}_{o})\big{]}=0$
@@
Given $\rho(A_{mm}-A_{mm}{(K_{m})}_{c}C_{mm}<1$, the quantity $\mathbb{E}[\lVert{(r_{m}^{t})}_{c}\rVert]$ is finite.
Therefore from inequality 46, we can say that $\mathbb{E}[\lVert{(\hat{h}_{m}^{t})}_{o}-{(\hat{h}_{m}^{t})}_{c}\rVert]\leq\delta_{max}^{m}$ where, $\delta_{max}^{m}$ is a function of $\mathbb{E}[\lVert{(r_{m}^{t})}_{c}\rVert],\mathbb{E}[\lVert{(r_{m}^{t})}_{o}\rVert],C_{mm},A_{mm},\&\sum_{n\neq m}A_{mn}{(\hat{h}_{n}^{t})}_{o}$.
@@
$\mathbb{E}[{(r_{m}^{t})}_{a}^{*}]=\mathbb{E}\Big{[}{y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{c}+A_{mm}\theta_{m}^{*}y_{m}^{t}\bigg{)}\Big{]}$ (47)
We also know from $\mathbb{E}\big{[}\nabla_{\hat{A}_{mn}}L_{s}\big{]}=0$ the following:
$\mathbb{E}\bigg{[}A_{mm}\theta_{m}^{*}y_{m}^{t}-\sum_{n\neq m}\hat{A}^{*}_{mn}{(\hat{h}_{n}^{t})}_{c}\bigg{]}=0$ (48)
@@
$\mathbb{E}[{(r_{m}^{t})}_{a}^{*}]=\mathbb{E}\bigg{[}{y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{c}+\sum_{n\neq m}\hat{A}^{*}_{mn}{(\hat{h}_{n}^{t})}_{c}\bigg{)}\bigg{]}$ (49)
@@
$\mathbb{E}[{(r_{m}^{t})}_{o}]=\bigg{[}{y_{m}^{t}}-C_{mm}\bigg{(}A_{mm}{(\hat{h}_{m}^{t})}_{o}+\sum_{n\neq m}{A}_{mn}{(\hat{h}_{n}^{t})}_{o}\bigg{)}\bigg{]}$ (50)
Subtracting equation 50 from 49, and using results from proposition 6.2 i.e., $\mathbb{E}[\lVert{(\hat{h}_{m}^{t})}_{o}-{(\hat{h}_{m}^{t})}_{c}\rVert]\leq\delta_{max}^{m}$ we obtain:
@@
Let $\sigma_{min}^{n}=\min_{n\neq m}\mathbb{E}[\lVert{(\hat{h}_{n}^{t-1})}_{o}\rVert]$, then using results of theorem 6.3 and inequality 52 we can infer that, $\big{\lVert}\sum_{n\neq m}[\hat{A}_{mn}^{*}-A_{mn}]\big{\rVert}\leq\frac{1}{\sigma_{min}^{n}}\cdot\bigg{(}\lVert A_{mm}\delta^{m}_{max}\rVert+\lVert\sum_{n,n\neq m}\hat{A}^{*}_{mn}\delta^{n}_{max}\rVert\bigg{)}$